# Occlusion-Aware Consistent Model Predictive Control for Robot Navigation in Occluded Obstacle-Dense Environments

### v4

@@ IV-C ADMM-based Optimization
where $G_{\text{obs},z}(\cdot)$ and $G_{\text{risk},z}(\cdot)$ represent inequality constraints for obstacle avoidance (11b) and risk boundary (11c), respectively. $H_{\text{kin},z}(\cdot)$ represents the equality constraint related to robot kinematics (11a). $\lambda_{\text{obs},z}$, $\lambda_{\text{risk},z}$, $\lambda_{\text{kin},z}$, $\lambda_{\text{cons},z}$ are the corresponding dual variables; $\rho_{\text{obs}}$, $\rho_{\text{risk}}$, $\rho_{\text{kin}}$, $\rho_{\text{cons}}$ are the corresponding penalty coefficients. Specifically, $\rho_{\text{cons}}$ is the penalty parameter for the proximal term of the consensus constraint. $\mathcal{I}_{+}(F)$ denotes an indicator function defined as follows:
@@
15:    Update ADMM dual variable $\lambda_{\text{cons},z}$ via (15b)
16:    for all sub-problems;
17:    Break if $\|\nabla L_{z}(\cdot)\|\leq 0.15$ and
18:    $\|s_{z}^{\iota+1}-\tilde{s}^{\iota+1}\|\leq 0.1$ and $\|\nabla\tilde{s}\|\leq 0.1$;
19:   End For
20:  Apply the first control input from the optimized
21:  consensus segment to the robot via (3);
22:End While
Note that the sub-problems are nearly independent with respect to different risk region configurations, except for the consensus constraint (10e). Meanwhile, the updating process of consensus variable $\tilde{s}$ depends on the results of all sub-problems. Therefore, the optimization problem can be decomposed into $N_{z}$ independent sub-problems. The consensus variable $\tilde{s}$ is updated once all sub-problems are optimized.
@@
For each sub-problem, (14a) is optimized using Newton’s method [35]. The dual variables corresponding to the inequality and equality constraints are updated in (14b)-(14d). The optimizations of all sub-problems are performed in parallel.
@@
$\|\nabla L_{z}(\cdot)\|\leq$ $\epsilon^{\text{dual}},\ \forall z\in\mathcal{I}_{0}^{N_{z}-1},$ (16a)
$\|s_{z}^{\iota+1}-\tilde{s}^{\iota+1}\|\leq$ $\xi^{\text{pri}},\ \forall z\in\mathcal{I}_{0}^{N_{z}-1},$ (16b)
@@
The optimization process continues until the termination criteria are met for all sub-problems and the consensus segment, or the maximum iteration $\iota_{\text{max}}=300$ is reached. A detailed description of the CMPC optimization process is provided in Algorithm 1.
@@
Figure 3: Snapshots of the simulation, where the robot navigates in an occluded, obstacle-dense environment. The robot, shown as an orange rectangle with black wheels, optimizes three trajectories in parallel. Dark-gray boxes represent visible obstacles, and light-gray boxes represent the occluded obstacles from the FoV of the robot. The arrow indicates the current velocity vector of the robot. Curves in different colors represent trajectories with different risk region configurations.
@@
TABLE II: Ablation Study Among Different Consensus Horizon Length
Consensus Horizon (s) Max Lat. Vel. Variance (m/s) Peak Lat. Acc. ($\text{m/s}^{2}$) Avg. Solving Time (ms)
@@
The computational complexity of the algorithms for each iteration can be specified as:
@@
  • Averaging operations over the consensus segment scale as $\mathcal{O}(N_{c}\cdot m)$, where $N_{c}$ is the consensus horizon and $m$ is the state dimension.
@@
Given that the planning horizon NN and consensus horizon $N_{c}$ are fixed, and the number of risk region $N_{\text{risk}}$ is bounded, the computational complexity per ADMM iteration scales linearly with the number of obstacles as $\mathcal{O}(N_{\text{obs}})$.
@@
In the Jacobi-Proximal ADMM-based optimization scheme, each sub-problem optimizes a single trajectory branch with its specific risk region configuration. In each iteration, the optimization step (14a) is solved in parallel for each risk region configuration $z$, yielding the optimal state sequence $s_{z}$ for each trajectory branch. This parallelization is key to achieving real-time performance. The solutions of all branches are then aggregated through the consensus update (15a)–(15b), which enforces that all branches share a common consensus segment, thus ensuring motion consistency. Additionally, the proximal term $\rho_{\text{cons}}|s_{z}-\tilde{s}|^{2}$ in the augmented Lagrangian (12) stabilizes the convergence of this parallel process, and this ensures numerical robustness.
@@ V Experiments
In this section, we present the experimental validation of our proposed CMPC for mobile robots navigating in occluded, obstacle-dense environments. The experiments are divided into two parts: simulations and real-world experiments.
@@ V-A1 Simulation Setup
The simulations are conducted using C++ and ROS1 on an Ubuntu 20.04 system, equipped with an Intel Ultra7 155H CPU (16 cores @ 1.40 GHz) and 16 GB of RAM. The simulation environment utilizes high-fidelity Gazebo 9 for dynamic simulation and RVIZ for visualization of the robot’s trajectories and motion. The robot’s dimensions are $800\,\text{mm}\times 400\,\text{mm}$. Blocks sized $1500\,\text{mm}\times 1500\,\text{mm}$ serve as both static and dynamic obstacles.
The robot is tasked with navigating through an occluded, obstacle-dense area containing both dynamic and static obstacles. Some dynamic obstacles are occluded from the robot’s FoV, such as the light-gray boxes inside the occluded region shown in Fig. 3(a) and Fig. 3(b). Dynamic obstacles initiate movement at a velocity of $0.6\,\text{m/s}$ to $1.0\,\text{m/s}$ along the lateral direction when their distance to the robot is within $2\,\text{m}$. The planning horizon is $N=6\,\text{s}$ with 44 steps per second, and the consensus segment is $N_{c}=2\,\text{s}$. The reference velocity of the robot is set to $1.8\,\text{m/s}$. The weighting matrices are set to $w_{\text{guide}}=3.5$, $w_{\text{vel}}=5.0$, $w_{\text{acc}}=1.8$.
@@
To further evaluate the impact of the consensus horizon, we perform an ablation study with different consensus horizons: $0\,\text{s}$ (no consensus), $1\,\text{s}$, $2\,\text{s}$, and $5\,\text{s}$ in simulation.
@@ V-A2 Results
Fig. 3 shows top-down view snapshots of the simulation using our CMPC. When the robot enters the occluded, obstacle-dense region at time instant $t=8\,\text{s}$, the dynamic obstacle outlined in the light-gray box is in the occlusion region blocked by Obs3, as shown in Fig. 3(a). When the robot continues to go forward, gets rid of the influence from Obs1 and approaches Obs3 at time instant $t=10\,\text{s}$, it starts to consider the risk regions behind Obs3, as shown in Fig. 3(b). The robot generates three trajectories: the light-blue trajectory neglects all the risk regions, which is the most aggressive one seeking the highest efficiency; the purple trajectory is medium aggressive considering the dynamic obstacle moving in medium velocity, and the dark-blue trajectory considers the worst-case where the dynamic obstacle moves in high velocity. When the dynamic obstacle is fully observable, as shown in Fig. 3(c), the robot can smoothly switch to the dark blue trajectory.
@@
Table I shows a performance comparison among different approaches. Notably, Control-Tree and the single hypothesis MPC without risk region considerations result in collisions with occluded obstacles, failing to ensure safety. A comparison of velocity profiles is shown in Fig. 4. While the single hypothesis MPC with risk region considerations avoids collisions, it exhibits a 38% higher lateral velocity variance and 51.7% higher peak lateral acceleration compared to our CMPC. The ablation study comparing CMPC-0 and our approach further demonstrates that the consensus segment effectively reduces lateral velocity variance by 15.3% and peak lateral acceleration by 47.9%, improving motion consistency significantly. Notably, the lateral velocity of these four baseline approaches increases significantly when encountering occluded dynamic obstacles, severely compromising motion consistency. These results indicate that the proposed CMPC strikes a better balance between conservatism and aggressiveness, ensuring both safety and improved motion consistency. Additionally, the average solving time of our CMPC is $40.78\,\text{ms}$, indicating that the CMPC can achieve real-time trajectory generation in obstacle-dense environments.
@@
The results in Table II reveal a critical trade-off when selecting the consensus horizon length. A consensus horizon of $2\,\text{s}$ achieves the best performance, yielding the lowest lateral velocity variance and peak lateral acceleration compared to the shorter consensus horizons of $0\,\text{s}$ and $1\,\text{s}$. However, extending the consensus horizon further to $5\,\text{s}$ leads to a performance degradation in these key metrics while increasing the average solving time. This decline in performance is attributed to the ’delayed decisions’ induced by an excessively long consensus horizon, which hinders the system’s ability to react agilely to changes in the environment. Therefore, future research will explore adaptive methods for selecting the optimal consensus horizon, dynamically adjusting $N_{c}$ based on environmental complexity and robot state to balance performance and efficiency.
@@
Figure 5: Snapshots of real-world experiments. The TianRacer robot successfully navigates through obstacles and avoids collision with the occluded dynamic obstacle.
@@ V-B1 Experiment Setup
The proposed CMPC strategy is deployed on a TianRacer robot, a mobile robot platform equipped with a four-wheel Ackermann-steering configuration, based on ROS1 platform. The robot’s dimensions are $380\,\text{mm}\times 210\,\text{mm}$. Static blocks sized $600\,\text{mm}\times 600\,\text{mm}$ serve as static obstacles, while dynamic obstacles are represented by other mobile robots moving at a constant velocity. The reference velocity of the robot is set to $0.5\,\text{m/s}$, and dynamic obstacles initiate movement at $0.2\,\text{m/s}$ when the robot is $0.5\,\text{m}$ away from them.
@@ VI Conclusions
In this study, we introduce a novel occlusion-aware CMPC for the safe navigation of mobile robots in occluded, obstacle-dense environments. The CMPC incorporates a module for modeling occluded regions and risk regions to proactively address potential safety threats from occluded obstacles. Coupled with a tree-structure CMPC and a common consensus segment, our strategy ensures safe traversal through occluded, obstacle-dense environments while maintaining motion consistency and task efficiency. Additionally, the integration of ADMM-based optimization enhances the computational efficiency of our strategy, enabling effective real-time trajectory generation. Extensive simulations demonstrate the effectiveness of our approach, achieving safe navigation with over 38% reduction in lateral velocity variance and over 51.7% lower peak lateral acceleration versus baselines. Real-world experiments on an Ackermann-steering mobile robot platform further demonstrate the effectiveness of our CMPC strategy, enabling safe navigation in physically occluded, obstacle-dense environments. As part of our future research, we will explore interactions with humans and robot agents in various physically occluded outdoor environments.

### v5

@@ IV-C ADMM-based Optimization
where $s_{z}$ is the robot’s state in the sub-problem with the $z$-th risk region configuration. $G_{\text{obs},z}(\cdot)$ and $G_{\text{risk},z}(\cdot)$ represent inequality constraints for obstacle avoidance (11c) and risk boundary (11d), respectively. $H_{\text{kin},z}(\cdot)$ represents the kinematic equality constraint (11b). $\lambda_{\text{obs},z}$, $\lambda_{\text{risk},z}$, $\lambda_{\text{kin},z}$, $\lambda_{\text{cons},z}$ are the corresponding dual variables; $\rho_{\text{obs}}$, $\rho_{\text{risk}}$, $\rho_{\text{kin}}$, $\rho_{\text{cons}}$ are the corresponding penalty coefficients. Specifically, $\rho_{\text{cons}}$ is the penalty parameter for the proximal term of the consensus constraint. $\mathcal{I}_{+}(F)$ is an indicator function defined as follows:
@@
15:    Update ADMM dual variable $\lambda_{\text{cons},z}$ via (15b);
16:    Break if the termination criteria (16a)-(16c)
17:    are met;
18:   End For
19:  Apply the first control input from the optimized
20:  consensus segment to the robot via (3);
21:End While
Note that the sub-problems are independent with respect to different risk region configurations, except for the consensus constraint (10e). Meanwhile, the updating process of consensus variable $\tilde{s}$ depends on the results of all sub-problems. Therefore, the optimization problem can be decomposed into $N_{z}$ independent sub-problems. The consensus variable $\tilde{s}$ is updated once all sub-problems are optimized.
@@
For each sub-problem, (14a) is optimized using Newton’s method [9]. The dual variables corresponding to the constraints are updated in (14b)-(14d). The optimizations of all sub-problems are performed in parallel.
@@
$\|\nabla L_{z}(\cdot)\|\leq$ $\epsilon^{\text{dual}},\quad\forall z\in\mathcal{I}_{0}^{N_{z}-1},$ (16a)
$\|s_{z}^{\iota+1}-\tilde{s}^{\iota+1}\|\leq$ $\xi^{\text{pri}},\quad\forall z\in\mathcal{I}_{0}^{N_{z}-1},$ (16b)
@@
The optimization process continues until the termination criteria (16a)-(16c) are met for all sub-problems and the consensus segment, or the maximum iteration $\iota_{\text{max}}=300$ is reached. A detailed description of the CMPC optimization process is provided in Algorithm 1.
@@
Figure 3: Simulation snapshots, where the robot navigates in an occluded, obstacle-dense environment. The robot, shown as an orange rectangle with black wheels, optimizes three trajectories in parallel. Dark-gray boxes represent visible obstacles, and light-gray boxes represent occluded obstacles. The arrow indicates the current velocity vector of the robot. Curves in different colors represent trajectories with different risk region configurations.
@@
TABLE II: Ablation Study Among Different Consensus Segment Length
Consensus Segment Length $T_{c}$ (s) Max Lat. Vel. Variance (m/s) Peak Lat. Acc. ($\text{m/s}^{2}$) Avg. Solving Time (ms)
@@
The computational complexity of the algorithm for each iteration can be specified as:
@@
  • Averaging operations over the consensus segment scale as $\mathcal{O}(N_{c}\cdot m)$.
@@
Given that the number of time steps $N$ in the planning horizon and the number of time steps $N_{c}$ in the consensus segment are fixed, and the number of risk regions $N_{\text{risk}}$ is bounded, the computational complexity per ADMM iteration scales with the number of obstacles as $\mathcal{O}(N_{\text{obs}})$.
@@
In the Jacobi-Proximal ADMM-based optimization scheme, each sub-problem optimizes a single trajectory branch with its specific risk region configuration. In each iteration, the optimization step (14a) is solved in parallel for each risk region configuration $z$, yielding the optimal states $s_{z}$ for each trajectory branch. This parallelization is key to achieving real-time performance. The solutions of all branches are then aggregated through the consensus update (15a)–(15b), which enforces that all branches share a common consensus segment, thus ensuring motion consistency. Additionally, the proximal term $\rho_{\text{cons}}\|s_{z}-\tilde{s}\|^{2}$ in the augmented Lagrangian (12) stabilizes the convergence of this parallel process, and this ensures numerical robustness.
@@ V-A1 Simulation Setup
The simulations are conducted using C++ and ROS1 on an Ubuntu 20.04 system, equipped with an Intel Ultra7 155H CPU (16 cores @ 1.40 GHz) and 16 GB of RAM. The simulation environment utilizes Gazebo 9 for dynamic simulation and RVIZ for visualization of the robot’s trajectories and motion. The robot’s dimensions are $800\,\text{mm}\times 400\,\text{mm}$. Blocks sized $1500\,\text{mm}\times 1500\,\text{mm}$ serve as both static and dynamic obstacles.
The robot is tasked with navigating through an occluded, obstacle-dense area containing dynamic and static obstacles. Some dynamic obstacles are occluded from the robot’s FoV, such as the light-gray boxes inside the occluded region shown in Fig. 3(a) and Fig. 3(b). Dynamic obstacles initiate movement at a velocity of $0.6\,\text{m/s}$ to $1.0\,\text{m/s}$ along the lateral direction when their distance to the robot is within $2\,\text{m}$. The planning horizon is $T_{h}=6\,\text{s}$ with $4$ steps per second, resulting in $N=24$ time steps; and the consensus segment length is $T_{c}=2\,\text{s}$, resulting in $N_{c}=8$ time steps. The reference velocity of the robot is set to $1.8\,\text{m/s}$. The weighting matrices are set to $w_{\text{guide}}=3.5$, $w_{\text{vel}}=5.0$, $w_{\text{acc}}=1.8$.
@@
To further evaluate the impact of the consensus segment, we perform an ablation study with different consensus segment lengths: $0\,\text{s}$ (no consensus), $1\,\text{s}$, $2\,\text{s}$, and $5\,\text{s}$.
@@ V-A2 Results
Fig. 3 shows top-down view snapshots of the simulation using our CMPC. When the robot enters the occluded, obstacle-dense region at time instant $t=8\,\text{s}$, a dynamic obstacle outlined in light-gray box is in the occlusion region blocked by Obs3, as shown in Fig. 3(a). When the robot continues to go forward and approaches Obs3 at time instant $t=10\,\text{s}$, it considers the risk regions behind Obs3, as shown in Fig. 3(b). The robot generates three trajectories: the dark-blue trajectory seeks the highest efficiency; the light-blue trajectory is medium aggressive, and the purple trajectory considers the worst-case to maximize safety awareness. When the dynamic obstacle is fully observable, as shown in Fig. 3(c), the robot can smoothly switch to the dark-blue trajectory.
@@
Table I shows a performance comparison among different approaches. Notably, Control-Tree and the single hypothesis MPC without risk region considerations result in collisions, failing to ensure safety. A comparison of velocity profiles is shown in Fig. 4. While the single hypothesis MPC with risk region considerations avoids collisions, it exhibits a 38% higher lateral velocity variance and 51.7% higher peak lateral acceleration compared to our CMPC. The ablation study comparing CMPC-0 and our approach further demonstrates that the consensus segment reduces lateral velocity variance by 15.3% and peak lateral acceleration by 47.9%, improving motion consistency significantly. Notably, the lateral velocity of these four baseline approaches increases significantly when encountering occluded dynamic obstacles, severely compromising motion consistency. These results indicate that our CMPC strikes a better balance between conservatism and aggressiveness, ensuring both safety and improved motion consistency. Additionally, the average solving time of our CMPC is $40.78\,\text{ms}$, indicating that it can achieve real-time trajectory generation in obstacle-dense environments.
@@
The results in Table II reveal a critical trade-off when selecting the consensus segment length. A consensus segment length of $2\,\text{s}$ achieves the best performance, yielding the lowest lateral velocity variance and peak lateral acceleration compared to the shorter consensus segment lengths of $0\,\text{s}$ and $1\,\text{s}$. However, extending the consensus segment length further to $5\,\text{s}$ leads to a performance degradation in these key metrics while increasing the average solving time. This decline in performance is attributed to the “delayed decisions” induced by an excessively long consensus segment, which hinders the system’s ability to react agilely to changes in the environment. Therefore, future research will explore adaptive methods for dynamically adjusting the consensus segment length based on environmental complexity and robot state to balance performance and efficiency.
@@
Figure 5: Real-world experiment snapshots on the TianRacer robot. The robot successfully navigates through obstacles and avoids collision with the occluded dynamic obstacle.
@@ V-B1 Experiment Setup
The proposed CMPC strategy is deployed on a TianRacer robot, a mobile robot platform equipped with a four-wheel Ackermann-steering configuration, based on ROS1. The robot’s dimensions are $380\,\text{mm}\times 210\,\text{mm}$. Static blocks sized $600\,\text{mm}\times 600\,\text{mm}$ serve as static obstacles, while dynamic obstacles are represented by other mobile robots. The reference velocity of the robot is set to $0.5\,\text{m/s}$, and dynamic obstacles initiate movement at $0.2\,\text{m/s}$ when the robot is $0.5\,\text{m}$ away from them.
@@ VI Conclusions
In this study, we introduce a novel occlusion-aware CMPC for the safe navigation of mobile robots in occluded, obstacle-dense environments. The CMPC incorporates a module for modeling occluded regions and risk regions to proactively address potential safety threats from occluded obstacles. Coupled with a tree-structure CMPC and a common consensus segment, our strategy ensures safe traversal through occluded, obstacle-dense environments while maintaining motion consistency and task efficiency. Additionally, the integration of ADMM-based optimization enhances the computational efficiency of our strategy, enabling real-time trajectory generation. Extensive simulations demonstrate the effectiveness of our approach, achieving safe navigation with 38% reduction in lateral velocity variance and 51.7% lower peak lateral acceleration versus baselines. Real-world experiments on an Ackermann-steering mobile robot platform further demonstrate the effectiveness of our CMPC strategy, enabling safe navigation in physically occluded, obstacle-dense environments. As part of our future research, we will investigate adaptive adjustment of the consensus segment length to further balance performance and efficiency.